### v1

@@ 1 Introduction
Another challenge in applying gradient-based methods to bilevel problems is the choice of step-size for the upper-level optimization. Convergence analysis for bilevel methods typically assumes highly accurate hypergradients and the use of a fixed, small step-size [23, 31, 32]. In the context of piggyback-style computed hypergradients, existing analyses also require highly accurate hypergradients and a sufficiently small, fixed step-size [28]. However, in practice, determining an appropriate step-size when working with approximate hypergradients is non-trivial. Assuming access to inexact lower-level solutions with controllable error and inexact hypergradients with a-posteriori error bounds, a condition was developed in [33] to adaptively select step-sizes and the required accuracy, ensuring convergence in a robust and cost-efficient manner. To connect and apply the method in [33] to primal-dual style differentiation, we propose an a-posteriori error bound analysis for the lower-level and adjoint solutions, as well as a-posteriori error bounds for the inexact hypergradient. We demonstrate the effectiveness of our approach through some bilevel learning tasks, including learning the discretization of the total variation (TV) and training input-convex neural networks (ICNNs) regularizers [34, 12]. The proposed method provides a framework that balances computational efficiency with rigorous error control, making it well-suited for large-scale bilevel learning applications.
@@ 2.1 Problem Setting
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\{h(x,y,K)\coloneqq\langle Kx,y\rangle+g(x)-f^{*}(y)\}.$ (S)
@@
${(\hat{x}(K),\hat{y}(K))}$ $\coloneqq\arg\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}h(x,y,K).$ (4b)
The upper-level loss function with respect to the lower-level solution $(\hat{x}(K),\hat{y}(K))$ is represented by $\ell$. Throughout the paper, we make the following assumption on $\ell$.
@@
Since the upper-level problem (4a) can be large-scale, we are interested in gradient-based bilevel methods [15, 23] which employ $\nabla\mathcal{L}(K)$ to solve (4a). To ensure a well-defined $\nabla\mathcal{L}(K)$ that can be approximated with theoretical guarantees, similar to [28], we consider the following regularity assumptions.
Assumption 2.
$g$ and $f^{*}$ are $\mu_{g}$- and $\mu_{f^{*}}$-strongly convex, respectively. Moreover, $g^{*}$ and $f$ are locally $C^{2,1}$. [^1] For simplicity, we consider locally $C^{2,1}$ functions, but the analysis can be generalized to locally $C^{2,\alpha}$ functions for $\alpha\in(0,1]$, similar to [28]
Under Assumption 1 and 2, $(\hat{x}(K),\hat{y}(K))$ given $K$ is unique, and the hypergradient can be calculated as follows:
@@
$+\langle\nabla\ell_{1}(\hat{x}(K)),X\rangle+\langle\nabla\ell_{2}(\hat{y}(K)),Y\rangle,$
@@
([37]) Let $\Phi:\mathcal{X}\rightarrow\mathbb{R}$ be $\mu$-strongly convex and differentiable. Denoting the minimizer of $\Phi$ by $x^{*}\in\mathcal{X}$, for all $x\in\mathcal{X}$ we have
@@
Let 2 hold. For a given operator $K$, and considering the unique saddle-point $(\hat{x},\hat{y})$ of (S), for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$ we have
@@
Since $\hat{x}$ is the minimizer of (P), and from 2 the primal problem’s objective $f(K\cdot)+g(\cdot)$ is $\mu_{g}$-strongly convex, utilizing Lemma 1 one can derive (6a). On the other hand, we note that the dual problem (D) is equivalent to :
@@ 2.2 Piggyback Algorithm
To compute the hypergradient $\nabla\mathcal{L}(K)$, two variations of the piggyback algorithm were introduced in [29]. By looking at (5), computing the hypergradient involves solving two saddle-point problems: (S) and (AS), respectively. A well-suited algorithm to solve each of these problems is the Primal-Dual hybrid gradient (PDHG) [36, Algorithm 3]. Since the adjoint problem (AS) depends on the solution of (S), two different approaches can be taken. One strategy is to solve (S) and (AS) alternatively, where each iterations includes a step of PDHG applied to (S), followed by plugging the result in (AS) and performing one step of inexact PDHG applied to the adjoint problem, as detailed in [38]. It was shown in [28, Theorem 2.2] that under Assumptions 1 and 2, and with a suitable choice of parameters for PDHG, this alternative piggyback algorithm converges linearly to $(\hat{x}(K),\hat{y}(K),\hat{X}(K),\hat{Y}(K))$, and thus to the hypergradient $\nabla\mathcal{L}(K)$.
@@
2:Set $\tau=\frac{\nu}{2\mu_{g}}$, $\sigma=\frac{\nu}{2\mu_{f^{*}}}$, and $\theta\in[\frac{1}{1+\nu},1]$.
3:Solve (S) using PDHG to find $(\tilde{x}(K),\tilde{y}(K))$ as an approximation of $(\hat{x}(K),\hat{y}(K))$.
4:Solve (ASI) using PDHG to find $(\tilde{X}(K),\tilde{Y}(K))$ as an approximation of $(\bar{X}(K),\bar{Y}(K))$
@@
([28, Theorem 2.3]) Under Assumption 1 and 2, let $\nu\leq\frac{2\sqrt{\mu_{g}\mu_{f^{*}}}}{\|K\|}$, $\tau=\frac{\nu}{2\mu_{g}}$, $\sigma=\frac{\nu}{2\mu_{f^{*}}}$, $\theta\in[\frac{1}{1+\nu},1]$ chosen as in [36, Algorithm 3]. The iterates of Algorithm 1 satisfy
@@ 3.1 Inexact Piggyback Algorithm and A-Posteriori Error Analysis
Let 2 hold. For a given operator $K$, and consider the unique saddle-point $(\bar{X},\bar{Y})$ of (ASI), we have that, for all $(X,Y)\in\mathcal{X}\times\mathcal{Y}$,
@@
Using 1, 2 and Lemma 1 we derive (9a) and (9b) as required. ∎
@@
where the last inequality is due to [39, Theorem 1] and the $L_{\nabla g}$ smoothness of $g$. Since $g^{*}$ is $C^{2,1}$ and $g$ is $L_{\nabla g}$ smooth, we have:
@@
Let Assumption 1 and 2 hold. Moreover, let $g$ and $f^{*}$ be $L_{\nabla g}$ and $L_{\nabla f^{*}}$ smooth, respectively. For $\epsilon^{x},\epsilon^{y},\delta^{X},\delta^{Y}>0$, if $\|\tilde{x}-\hat{x}\|\leq\epsilon^{x}$, $\|\tilde{y}-\hat{y}\|\leq\epsilon^{y}$. Let $(\tilde{X},\tilde{Y})\in\mathcal{X}\times\mathcal{Y}$ such that
@@
where the last inequality again utilizes [39, Theorem 1]. Moreover, we have:
@@
Now, utilizing lemma 1 and theorem 2, we propose an inexact variation of Piggyback algorithm with controllable error on the lower-level iterates as well as the hypergradient, which is outlined in Algorithm 2.
@@ 3.2 Method of Adaptive Inexact Descent (MAID)
Given the approximate hypergradient provided by Algorithm 2, we employ the adaptive backtracking line search scheme introduced in [33] to find suitable upper-level step-sizes, and to adaptively adjust the tolerances $\epsilon^{x},\epsilon^{y},\delta^{X},\delta^{Y}$ required for calculating an approximate hypergradient. The following lemma gives a sufficient decrease condition which we will deploy as the line search rule for upper-level iterations.
@@
[33, Lemma 3.4] Let 1 hold, $\lambda\in\mathbb{R}$, and suppose that the upper-level loss $\ell$ is convex. For each upper-level iteration $t=0,1,\dots$, set $u_{t}\coloneqq(\tilde{x}(K_{t}),\tilde{y}(K_{t}))$, $\bar{\epsilon}=\max\{\epsilon^{x}_{t},\epsilon^{x}_{t+1},\epsilon^{y}_{t},\epsilon^{y}_{t+1}\}$, and denoting
@@
We denote the $t$-th iterate of the linear learnable parameters by $K_{t}$. Considering the bilevel problem (4) and utilizing lemma 5, the adaptive backtracking line search with restart, outlined in algorithm 3, is employed to solve the upper-level problem (4a).
Algorithm 3 [33, Algorithm 3.1] Method of Adaptive Inexact Descent (MAID). Hyperparameters: $\underline{\rho},\overline{\rho}\in(0,1)$ control the reduction and increase of the step-size $\alpha_{k}$, respectively; $\underline{\nu},\overline{\nu}\in(0,1)$ govern the reduction and increase of accuracies $\delta_{k}$ and $\epsilon_{k}$; $\max_{\text{BT}}\in\mathbb{N}$ is the maximum number of backtracking iterations.
@@ 4.1 Learning the Discretization of Total Variation
In the work of [40], the need for a new discretization for the total variation was identified. The aim was to enforce common pixel grid locations of the employed vector fields which in turn promotes rotational invariance that is not granted with the standard discretization. This was later extended to a general (consistent) setting using learnable interpolation filters for the total variation and its second-order extension [29, 41]. This is treated in the setting of standard linear inverse problems, where an image $x\in\mathbb{R}^{M\times N}$ is sought to be reconstructed. As usual, we are given a corrupted observation $u$ and the inverse problem is governed by a $\mu_{g}$-strongly convex data fidelity term $g(x,u)$ with a linear operator $A$.
@@
Combining (28) with a convex data fidelity term $g(x,u)$ in an inverse problem setting, this yields the subsequent saddle-point problem with regularisation parameter $\lambda\in\mathbb{R}^{+}$:
$\min_{x,q}\max_{y}g(x,u)+\lambda\|q\|_{Z}+\langle Dx-K^{*}q,y\rangle.$ (29)
@@
$\min_{X,Q}\max_{Y}\langle DX-K^{*}Q,Y\rangle+\tfrac{1}{2}\langle\nabla^{2}g(\hat{x},u)X,X\rangle\\ +\tfrac{1}{2}\langle\nabla^{2}\lambda\|\hat{q}\|_{Z},Q\rangle+\langle\nabla\ell(\hat{x},x^{*}),X\rangle.$ (30)
Note that in (29), the non-smooth absolute function $\|\cdot\|_{2,1,1}$ does not fully satisfy the regularity assumptions on $g$. While this would typically require smoothing, we observe that the primal-dual style differentiation approach in bilevel learning performs robustly even in this less regular setting, similar to the findings in [28]. Therefore, we directly work with (29).
For a given set of corresponding ground truth images and corrupted observations $(x_{i}^{*},u_{i})_{i}$, the standard bilevel setting in (4) can be employed, which – given the saddle-point structure of the underlying lower-level problem – can be approached using Algorithm 3. The experiments were conducted for the inverse problem of image denoising for a randomly sampled train subset of the BSDS train data set [42] with 32 images of size $128\times 128$. To generate the corrupted observations for the image denoising task, zero-mean additive Gaussian noise $\sim\mathcal{N}(0,\sigma^{2})$ with $\sigma=25.5$ was used.
@@
The linear operator $K$ was heuristically chosen to consist of $n=8$ filters with filter kernels $5\times 5$. The parameters in Algorithm 3 were set to $\underline{\nu}=0.5$, $\bar{\nu}=1.05$, $\underline{\rho}=0.5$, $\bar{\rho}=\tfrac{10}{9}$ following [33]. For comparison between the adaptive and non-adaptive settings, where the latter corresponding to the native inexact version of piggyback, the initial values for $\epsilon_{0}$, $\delta_{0}$ and $\alpha_{0}$ were kept the same across both cases. However, in the non-adaptive version, these parameters were fixed throughout the algorithm. Figure 1(a) shows the learned filters with the adaptive Algorithm 3 and without for $\epsilon_{0}=10^{-3}$, $\delta_{0}=10^{-3}$ and $\alpha_{0}=10^{-2}$, showing that both settings seem to capture very related orientations. While clearly the quantitatively learned filters are very similar, it is of higher interest to investigate the resulting implications on the used computational budget in terms of required lower-level iterations. Therefore, Figure 1(b) shows the resulting upper-level loss depending on the required total computational budget in all standard and adjoint primal-dual algorithms for different experimental settings. The benefit of adaptively choosing the step-size can clearly be seen as the resulting upper-level loss decreases smoothly with increasing computational budget in the lower-level solver. On the other hand, the dashed curves show the scenarios with the same starting accuracies and learning rate $\alpha$ as in the adaptive setting, however, these parameters were kept constant during the reconstruction algorithm, where PDHG was adapted for the saddle-point problem in (29). Depending on the initial learning rate, the upper-level loss even starts oscillating after an initial phase of decrease.

### v2

@@ 1 Introduction
A class of methods for solving bilevel problems involves reducing them to single-level formulations. One such approach, proposed in [31, 32], utilizes a single step of a primal-dual algorithm to approximate the lower-level solution. The adjoint problem, arising from the IFT approach, is then solved using either Conjugate Gradient (CG) with high accuracy or an analytic exact inversion of the lower-level Hessian in [31], while [32] employs a general linear solver. Although these methods do not require a highly accurate lower-level solution at each iteration, they do not provide adaptivity in determining the necessary accuracy for solving the lower-level problem to ensure progress in the upper-level iterations. This results in a high number of upper-level iterations, and their global convergence relies on assumptions regarding initialisation.
Another group of single-level reduction methods, known as fully first-order methods [33, 34, 35], replaces the lower-level problem with a so-called value function [36]. These methods reformulate the bilevel problem as a constrained optimization problem, avoiding the need for IFT-based evaluations and thus eliminating the computational burden of computing or approximating the inverse Hessian of the lower-level problem. Instead, they approximate the lower-level solution using a fixed number of iterations across all upper-level iterations. However, they require careful tuning of step-sizes and penalty multipliers to ensure convergence. Notably, the majority of works employing these approaches focus on stochastic bilevel problems.
@@
Another challenge in applying gradient-based methods to bilevel problems is the choice of step-size for the upper-level optimization. Convergence analysis for nested gradient-based bilevel methods typically assumes highly accurate hypergradients and the use of a fixed, small step-size [23, 37, 38]. In the context of piggyback-style computed hypergradients, existing analyses also require highly accurate hypergradients and a sufficiently small, fixed step-size [28]. However, in practice, determining an appropriate step-size when working with approximate hypergradients is non-trivial. Assuming access to inexact lower-level solutions with controllable error and inexact hypergradients with a-posteriori error bounds, a condition was developed in [39] to adaptively select step-sizes and the required accuracy, ensuring convergence in a robust and cost-efficient manner. To connect and apply the method in [39] to primal-dual style differentiation, we propose an a-posteriori error bound analysis for the lower-level and adjoint solutions, as well as a-posteriori error bounds for the inexact hypergradient. We demonstrate the effectiveness of our approach through some bilevel learning tasks, including learning the discretization of the total variation (TV) and training input-convex neural networks (ICNNs) regularizers [40, 12]. The proposed method provides a framework that balances computational efficiency with rigorous error control, making it well-suited for large-scale bilevel learning applications.
To summarize, our contributions are as follows:
  • We provide a detailed theoretical analysis establishing a-posteriori controllable error bounds for primal-dual-style bilevel differentiation, extending the framework in [39].
  • We prove a convergence result for primal-dual-style bilevel methods, which, to the best of our knowledge, had not been established before.
  • We demonstrate the suitability of our method to learn improved discretizations of the Total Variation (TV) and to train input convex neural networks (ICNNs) as regularizers via bilevel learning. In the latter case, we show improved performance over previous methods for training such data-adaptive regularizers.
@@ 2.1 Problem Setting
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\{h(x,y;K)\coloneqq\langle Kx,y\rangle+g(x)-f^{*}(y)\}.$ (S)
@@
${(\hat{x}(K),\hat{y}(K))}$ $\coloneqq\arg\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}h(x,y;K).$ (4b)
To define the upper-level loss with respect to the lower-level solution, we assume that $(\hat{x}(K),\hat{y}(K))$ given $K$ is unique. This can be guaranteed if $f^{*},g$ are strongly convex.
Assumption 1.
$f^{*}$ and $g$ are $\mu_{f^{*}}$- and $\mu_{g}$-strongly convex, respectively.
Remark 1.
In particular, this implies that $f$, $g^{*}$ have Lipschitz gradient, i.e. $f$, $g^{*}$ are $C^{1,1}$ [43].
Here, the upper-level loss function with respect to the lower-level solution is represented by $\ell$. Throughout the paper, we make the following assumption on $\ell$.
@@
Since the upper-level problem (4a) can be large-scale, we are interested in gradient-based bilevel methods [15, 23] which employ $\nabla\mathcal{L}(K)$ to solve (4a). Under 1 and 2, $(\hat{x}(K),\hat{y}(K))$ given $K$ is unique, and the hypergradient can be calculated as follows [28]:
@@
$+\langle\nabla\ell_{1}(\hat{x}(K)),X\rangle+\langle\nabla\ell_{2}(\hat{y}(K)),Y\rangle.$
Given a convex function $g$, Alexandrov’s theorem [44] states that the set-valued subgradient map $\partial g$ is differentiable almost everywhere. Moreover, at a point of differentiability $x$, $\partial g(x)$ is single valued and hence $\nabla g(x)$ exists. We then say that $\nabla^{2}g(x)$ exists and is equal to $\nabla(\nabla g)(x)$. In particular, $\nabla^{2}g(x)$ is non-singular [28] if $g$ is also strongly convex. In (AS) and what follows, both $\nabla^{2}g$ and $\nabla^{2}f^{*}$ will be interpreted in this sense.
@@
([45]) Let $\Phi:\mathcal{X}\rightarrow\mathbb{R}$ be $\mu$-strongly convex and differentiable [^1] The above lemma still holds when $\Phi$ is just $\mu$-strongly convex. The gradient $\nabla\Phi$ can be replaced by any subgradient $g_{x}\in\partial\Phi(x)$ [46, Theorem 5.24(iii)].. Denoting the minimizer of $\Phi$ by $x^{*}\in\mathcal{X}$, for all $x\in\mathcal{X}$ we have
@@
Let 1 hold. For a given operator $K$, and considering the unique saddle-point $(\hat{x},\hat{y})$ of (S), for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$ we have [^2] By slight abuse of notation, we still denote any subgradients with $\nabla g(x)$ and $\nabla f^{*}(y)$.
@@
Since $\hat{x}$ is the minimizer of (P), and from 1 the primal problem’s objective $f(K\cdot)+g(\cdot)$ is $\mu_{g}$-strongly convex, utilizing Lemma 1 one can derive (6a). On the other hand, we note that the dual problem (D) is equivalent to :
@@ 2.2 Piggyback Algorithm
To compute the hypergradient $\nabla\mathcal{L}(K)$, two variations of the piggyback algorithm were introduced in [29]. By looking at (5), computing the hypergradient involves solving two saddle-point problems: (S) and (AS), respectively. A well-suited algorithm to solve each of these problems is the Primal-Dual hybrid gradient (PDHG [^3] The authors in [47] referred to this algorithm as the modified primal-dual hybrid gradient (PDHGM).) [42, Algorithm 3]. Since the adjoint problem (AS) depends on the solution of (S), two different approaches can be taken. One strategy is to solve (S) and (AS) alternatively, where each iterations includes a step of PDHG applied to (S), followed by plugging the result in (AS) and performing one step of inexact PDHG applied to the adjoint problem, as detailed in [48].
@@
2:Set step-sizes $\tau=\frac{\nu}{2\mu_{g}}$, $\sigma=\frac{\nu}{2\mu_{f^{*}}}$, and extrapolation factor $\theta\in[\frac{1}{1+\nu},1]$.
3:Apply [42, Algorithm 3] to (S) with $\tau,\sigma,\theta$ chosen above to find approximate solution $(\tilde{x}(K),\tilde{y}(K))$.
4:Apply [42, Algorithm 3] to (ASI) with $\tau,\sigma,\theta$ chosen above to find approximate solution $(\tilde{X}(K),\tilde{Y}(K))$.
@@
To ensure a well-defined hypergradient that can be approximated with theoretical guarantees, similar to [28], we impose additional regularity assumption on $g^{*}$ and $f$.
Definition 1.
A $C^{2}$ function $f$ is said to be $C^{2,\alpha}$ if for some $L_{\nabla^{2}f}>0$, the following holds for all $x_{1},x_{2}$:
$\|\nabla^{2}f(x_{1})-\nabla^{2}f(x_{2})\|\leq L_{\nabla^{2}f}\|x_{1}-x_{2}\|^{\alpha}$ (9)
Assumption 3.
$f$ and $g^{*}$ are locally $C^{2,1}$. [^4] For simplicity, we consider locally $C^{2,1}$ functions, but the analysis can be generalized to locally $C^{2,\alpha}$ functions for $\alpha\in(0,1]$, similar to [28].
It was shown in [28, Theorem 2.2] that under 1, 2, and 3, and with a suitable choice of parameters for PDHG, i.e. step-sizes $\tau,\sigma$ and extrapolation factor $\theta$, this alternative piggyback algorithm converges linearly to $(\hat{x}(K),\hat{y}(K),\hat{X}(K),\hat{Y}(K))$, and thus to the hypergradient $\nabla\mathcal{L}(K)$.
@@
([28, Theorem 2.3]) Under 1, 2, and 3, let $\nu\leq\frac{2\sqrt{\mu_{g}\mu_{f^{*}}}}{\|K\|}$, $\tau=\frac{\nu}{2\mu_{g}}$, $\sigma=\frac{\nu}{2\mu_{f^{*}}}$, $\theta\in[\frac{1}{1+\nu},1]$ chosen as in [42, Algorithm 3]. The iterates of Algorithm 1 satisfy
@@ 3.1 Inexact Piggyback Algorithm and A-Posteriori Error Analysis
Let 1 and 2 hold. For a given operator $K$, and consider the unique saddle-point $(\bar{X},\bar{Y})$ of (ASI), we have that, for all $(X,Y)\in\mathcal{X}\times\mathcal{Y}$,
@@
Using 1, 2, and Lemma 1 we derive (10a) and (10b) as required. ∎
@@
where the last inequality is due to [43, Theorem 1] and the $L_{\nabla g}$ smoothness of $g$. Since $g^{*}$ is $C^{2,1}$ and $g$ is $L_{\nabla g}$ smooth, we have:
@@
Suppose 1, 2, and 3 hold. Moreover, let $g$ and $f^{*}$ be $L_{\nabla g}$ and $L_{\nabla f^{*}}$ smooth, respectively. For $\epsilon^{x},\epsilon^{y},\delta^{X},\delta^{Y}>0$, if $\|\tilde{x}-\hat{x}\|\leq\epsilon^{x}$, $\|\tilde{y}-\hat{y}\|\leq\epsilon^{y}$. Let $(\tilde{X},\tilde{Y})\in\mathcal{X}\times\mathcal{Y}$ such that
@@
where the last inequality again utilizes [43, Theorem 1]. Moreover, we have:
@@
Note that Theorem 2, and in particular (17), provides an a-posteriori and fully computable error bound for the inexact hypergradient calculated through primal-dual style differentiation, which can be used in adaptive bilevel methods such as [39]. Now, utilizing Lemma 1 and Theorem 2, we propose an inexact variation of Piggyback algorithm with controllable error on the lower-level iterates as well as the hypergradient, which is outlined in Algorithm 2.
@@ 3.2 Method of Adaptive Inexact Descent (MAID)
Given the approximate hypergradient provided by Algorithm 2, we employ the adaptive backtracking line search scheme introduced in [39] to find suitable upper-level step-sizes, and to adaptively adjust the tolerances $\epsilon^{x},\epsilon^{y},\delta^{X},\delta^{Y}$ required for calculating an approximate hypergradient. We denote the $t$-th iterate of the linear learnable parameters by $K_{t}$. The following lemma gives a sufficient decrease condition which we will deploy as the line search rule for upper-level iterations.
@@
[39, Lemma 3.5] Let 2 hold, $\lambda\in\mathbb{R}$, and suppose that the upper-level loss $\ell$ is convex. For each upper-level iteration $t=0,1,\dots$, set $u_{t}\coloneqq(\tilde{x}(K_{t}),\tilde{y}(K_{t}))$, $\bar{\epsilon}=\max\{\epsilon^{x}_{t},\epsilon^{x}_{t+1},\epsilon^{y}_{t},\epsilon^{y}_{t+1}\}$, and denoting
@@
Considering the bilevel problem (4) and utilizing Lemma 5, the adaptive backtracking line search with restart, outlined in Algorithm 3, is employed to solve the upper-level problem (4a). Note that, as detailed in the steps of Algorithm 3, Lemma 5 serves as an accuracy-dependent line search condition to adjust the required accuracy for progressing in the upper-level optimization iterations. When this condition is satisfied, it indicates that the accuracy is adequate for a sufficient decrease and suggests that the lower-level tolerance can be enlarged to reduce the computational cost of the next upper-level iteration. Conversely, if the condition does not hold and the maximum number of line search steps is reached, it indicates that the accuracy was insufficient. In this case, we proceed by shrinking the lower-level tolerance and restarting the computation of the lower-level solution, as well as rechecking the line search condition.
Given that Lemma 3 and Theorem 2 provide a practical and computable framework, as outlined in Algorithm 2, for computing an inexact hypergradient required in Algorithm 3, and that this aligns with the regularity assumptions needed in Lemma 5 as well as the convergence requirements of Algorithm 3, the convergence of our bilevel scheme follows from the theorem below.
Theorem 3.
[39, Theorem 3.19] Under 2 and 3, the iterates $K_{t}$ of Algorithm 3 satisfy
$\lim_{t\to\infty}\|\nabla\mathcal{L}(K_{t})\|=0.$
Algorithm 3 [39, Algorithm 3.1] Method of Adaptive Inexact Descent (MAID). Hyperparameters: $\underline{\rho}\in(0,1)$ and $\overline{\rho}>1$ control the reduction and increase of the step-size $\alpha_{k}$, respectively; $\underline{\nu}\in(0,1)$ and $\overline{\nu}>1$ govern the reduction and increase of accuracies $\delta_{k}$ and $\epsilon_{k}$; $\max_{\text{BT}}\in\mathbb{N}$ is the maximum number of backtracking iterations.
@@ 4.1 Learning the Discretization of Total Variation
In the work of [49], the need for a new discretization for the total variation was identified. The aim was to enforce common pixel grid locations of the employed vector fields which in turn promotes rotational invariance that is not granted with the standard discretization. This was later extended to a general (consistent) setting using learnable interpolation filters for the total variation and its second-order extension [29, 50]. This is treated in the setting of standard linear inverse problems, where an image $x\in\mathbb{R}^{M\times N}$ is sought to be reconstructed. In this experiment, we consider the task of image denoising. Given a noisy observation $u$, we consider the 1-strongly convex $L^{2}$ data fidelity $\tfrac{1}{2}\|x-u\|^{2}_{2}$.
@@
Combining (29) with the $L^{2}$ data fidelity, this yields the subsequent saddle-point problem with regularization parameter $\lambda\in\mathbb{R}^{+}$:
$\min_{x,q}\max_{y}\tfrac{1}{2}\|x-y\|^{2}_{2}+\lambda\|q\|_{Z}+\langle Dx-K^{*}q,y\rangle.$ (30)
@@
$\min_{X,Q}\max_{Y}\langle DX-K^{*}Q,Y\rangle+\tfrac{1}{2}\langle X,X\rangle\\ +\tfrac{1}{2}\langle\nabla^{2}\lambda\|\hat{q}\|_{Z},Q\rangle+\langle\nabla\ell(\hat{x},x^{*}),X\rangle.$ (31)
In (30), both the non-smooth absolute function $\|\cdot\|_{2,1,1}$ and the $L^{2}$ data term jointly constitute the original function $g$ defined in the primal objective in (P) of the lower-level problem. Note, however, that the non-smooth term $\|\cdot\|_{2,1,1}$ does not fully satisfy the regularity assumptions imposed on $g$ in (P). While this would typically require smoothing, we observe that the primal-dual style differentiation approach in bilevel learning performs robustly even in this less regular setting, similar to the findings in [28]. Therefore, we directly work with (30).
For a given set of corresponding ground truth images and corrupted observations $(x_{i}^{*},u_{i})_{i}$, the standard bilevel setting in (4) can be employed, which given the saddle-point structure of the underlying lower-level problem can be approached using Algorithm 3. The experiments were conducted for the inverse problem of image denoising for a randomly sampled train subset of the BSDS train data set [51] with 32 images of size $128\times 128$. To generate the corrupted observations for the image denoising task, zero-mean additive Gaussian noise $\sim\mathcal{N}(0,\sigma^{2})$ with $\sigma=25.5$ was used.
@@
The linear operator $K$ was heuristically chosen to consist of $n=8$ filters with filter kernels $5\times 5$. The parameters in Algorithm 3 were set to $\underline{\nu}=0.5$, $\bar{\nu}=1.05$, $\underline{\rho}=0.5$, $\bar{\rho}=\tfrac{10}{9}$ following [39]. For comparison between the adaptive and non-adaptive settings, where the latter corresponding to the native inexact version of piggyback, the initial values for $\epsilon_{0}$, $\delta_{0}$ and $\alpha_{0}$ were kept the same across both cases. However, in the non-adaptive version, these parameters were fixed throughout the algorithm. Figure 1(a) shows the learned filters with the adaptive Algorithm 3 and without for $\epsilon_{0}=10^{-3}$, $\delta_{0}=10^{-3}$ and $\alpha_{0}=10^{-2}$, showing that both settings seem to capture very related orientations. While clearly the quantitatively learned filters are very similar, it is of higher interest to investigate the resulting implications on the used computational budget in terms of required lower-level iterations. Therefore, Figure 1(b) shows the resulting upper-level loss depending on the required total computational budget in all standard and adjoint primal-dual algorithms for different experimental settings. The benefit of adaptively choosing the step-size can clearly be seen as the resulting upper-level loss decreases smoothly with increasing computational budget in the lower-level solver. On the other hand, the dashed curves show scenarios with the same initial accuracies and learning rate $\alpha$ as in the adaptive setting, but these parameters are now kept constant throughout the learning algorithm. Depending on the initial learning rate, the upper-level loss even starts oscillating after an initial phase of decrease.